Assume the risk free rate is 2.0%. The SP500 is considered the market. Today (T-0), you invest $400 in Stock A and $600 in Stock B to create Portfolio A,B. Assume there are not taxes or dividends. The one year performance of the stocks and the market is summarized in the table below. Use this information to help answer questions 16-20. Total Return 12.0% 18.0% 10.0% Total Risk 14,0% 15.0% 12.0% Investment Beta Market Stock A Stock B 1.00 400 600 0.60 16. Portfolio A,Bs beta is closest to: a. 0.80 b. 0.92 c. 1.00 d. 1.08 e. 1.20

17. If you doubled the size of your investment today (T=0), your total dollar return would be closest to:

a. $66 b. $92 c. $184 d. $264 e. $280
Answer #1

Answer:

16) option B:0.92

Stock A investment A=$400

Stock B investment B=$600

Beta of Stock  A Ba=1.4

Beta of Stock B Bb=0.60

Portfolio Beta = A/(A+B)*Ba+ B/(A+B)*Bb=0.4*1.4+0.6*0.60=0.92

17) Option D:$264

Since investment is doubled

so new investment in A=$800

Return from A=800*% return =800*18%=$144

New investment in B=$1200

Return from B=1200*% return =1200*10%=$120

total return=144+120=$264